true or false?



 A Suppose the consolidated balance sheet of an economy where the public holds all its money in the form of bank deposits is shown in the following table. If the banking system is originally in equilibrium and then the economy’s central bank sells 50 worth of government bonds to the banking system, immediately after the transaction there is no change in the money supply, but after the banking system has returned to equilibrium, the money supply is reduced by 1000. [Hint: Use two additional balance sheets in your answer.] Assets: 

Liabilities: 

Reserves 

100 

Deposits 

2000 

Government Bonds 

300 

Loans Outstanding 

1800 

Capital 

200 

Total 

2200 

Total 

2200 


1
Expert's answer
2021-04-05T07:22:24-0400

If the economy’s central bank sells 50 worth of government bonds to the banking system, and after the banking system has returned to equilibrium, the money supply is reduced by 1000, then it correlates with the data, that reserves are 100, deposits are 2000, so the reserve ratio is rr = 100/2000 = 0.05, and the money multiplier is mm = 1/rr = 1/0.05 = 20. As a result the money supply decreased by 1000, which is 20 times more than the 50 worth of government bonds.


So, the statement is true.